Find the determinant of this 3x3 matrix using expansion by minors about the second row
0 -4 6
A= 5 0 0
9 -2 7

Answer by Alan3354(69443)

Find the determinant of this 3x3 matrix using expansion by minors about the second row
0 -4 6
A= 5 0 0
9 -2 7
============
= -5*(-4*7 - (-2*6)) = -5*(-28 + 12) = 80 (the other minors are zeroes)